A battery is connected in series with a variable resistor. When the resistor is 8.50 ȍ, the current is 1.00 A and when the resistor is 6.50 ȍ, the current is 1.25 A. What is the value of the internal resistance?


1
Expert's answer
2021-12-27T08:20:58-0500

Answer

Given data

First case

Resistance R1=8.5"\Omega"

Current I1=1A

Now second case

Resistance R2=6.50"\Omega"

Current I2=1.25A

Now using the ohms law formula

"I=\frac{E}{r+R}"

Where r us internal resistance

Now write equation for both case

"1=\frac{E}{r+8.5}.........(1)"

"1.25=\frac{E}{r+6.50}.........(2)"

Divided equation 1 by 2 and solve this

We get internal resistance

r=1.5"\Omega"